OpenStudy (anonymous):

Evaluate the following expression for r = 5 and q = -6 5r – (3r + 2q) + 4(2 – q)

OpenStudy (vishweshshrimali5):

for r = 5 and q = -6 5r - (3r + 2q) +4(2-q) = 5*5 - (3*5 + 2 *(-6) ) + 4(2-(-6)) = 25 - (15 -12) + 4*8 = 25 - 3 + 32 = 54 this is your answer

OpenStudy (aroub):

just substitute r with 5 and q with -6 and the simplify thats all !! :)

OpenStudy (vishweshshrimali5):

hey am i right ?

OpenStudy (aroub):

wait ill do ot give me 2 min

OpenStudy (anonymous):

Evaluate the following expression for a = -4 and c = 3 5a – 4c + 6(a – 5c) – 2(a + 3)

OpenStudy (aroub):

25 -15-12+8+6 =12 i got 12

OpenStudy (aroub):

^^ @vish...

OpenStudy (vishweshshrimali5):

sorry aroub but this time also i believe that i am right

OpenStudy (aroub):

the second question : -20-12-24-15+8-6=

OpenStudy (vishweshshrimali5):

-69

OpenStudy (aroub):

u know i think ur right i should have put instead of -12 , +12 im really so sorry

OpenStudy (aroub):

yes 54 is correct